5. A call option on Company B common stock is worth $8 with 7 months before expiration. The strike price on the call is $40 and the price per share is currently trading at $44 per share. The put option at the same exercise price is worth $1.50.

a. Is the call option in or out or the money?

b. Is the put option in or out of the money?

c. At what extra above expiration value is the call selling for?

d. At what extra above expiration value is the put selling for?

e. If you bought both the put and call in (a) above and held them both to expiration, calculate your loss or profit on both the put and call if the price at expiration was $47 per share.

f. Graph (contingency graph) the payoffs for a long straddle at expiration [stock price X axis, option payoff on the Y axis] at different stock prices at the expiration of a straddle. X=40     Call = 8   Put = 1.50
Answer #1

1.
Call is in the money

2.
Put is out of the money

3.
=8-MAX(44-40,0)=4

4.
=1.50-MAX(40-44,0)=1.50

P.S.: I am not allowed to answer more than 4 questions